I need help. The directions are solve each system. If the system has in infinate solution set, specify it and give three solutions. If the system has no solution, say so.
The prolem is:
d = 2 - 6c
(1/2)d - c = 1

Answer by stanbon(75887)

The prolem is:
d = 2 - 6c
(1/2)d - c = 1
-----
Substitute for "d" and solve for "c":
(1/2)(2-6c) - c = 1
----
1 - 3c - c = 1
-4c = 0
c = 0
--------
Solve for "d" using d = 2 - 6c
d = 2 - 6*0
d = 2
===============
Cheers,
Stan H.
===============